OpenStudy (anonymous):

explain how to find the equation of the line, in standard form and slope–intercept form, passing through (–1, –7) and (1, –1). Also compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope–intercept form.

OpenStudy (katrinakaif):

Do you have any ideas first?

OpenStudy (katrinakaif):

To find an equation, our first and main step is finding the slope of the two given points To do so, we use the equation : y2-y1 ----- x2-x1 Your given points: (-1,-7) : (x1,y1) (1,-1) : (x2,y2)

OpenStudy (anonymous):

ok

OpenStudy (katrinakaif):

Then we use the equation: (y-y1) = m (x-x1) Subsituting m with what we concluded with in Step 1 and picking any one pair of points to substitute as (x1,y1)

OpenStudy (anonymous):

ok then?

OpenStudy (katrinakaif):

Lol You get your equation.

OpenStudy (anonymous):

k but what are the benefits of writing an equation in standard form or slope intercept form????

OpenStudy (katrinakaif):

What do you think?

OpenStudy (anonymous):

you just showed me how to do the slope intercept form right? how do i do it in standard form?????????

OpenStudy (anonymous):

????

OpenStudy (katrinakaif):

You try.. Standard equation is simply: ax + by = c Use this to guide you: http://jwilson.coe.uga.edu/emt668/emat6680.2002/jackson/chapter%205%20lesson%20plan/day6.html

OpenStudy (anonymous):

ok the slope intercept is 0 correct?

OpenStudy (katrinakaif):

y2-y1 ----- x2-x1 Your given points: (-1,-7) : (x1,y1) (1,-1) : (x2,y2) -1 - (-7) -------- 1 - (-1) I dont believe this gives you 0. Try again.

OpenStudy (anonymous):

3

OpenStudy (katrinakaif):

Try once again. :/

OpenStudy (anonymous):

-1-(-7)=6 1-(-1)=2 6/2=3

OpenStudy (katrinakaif):

Ah you are right! Apologies :] Now use that slope and plug it into the formula: y-y1 = m (x-x1) while m = 3 and (x1,y1) is equal to any given pair of points you are given

OpenStudy (anonymous):

i still dont know how to figure out the standard form i know slope just need help in the other

OpenStudy (katrinakaif):

Well lets start with the slope intercept form, What did you get?

OpenStudy (anonymous):

i got 3

OpenStudy (katrinakaif):

That is the slope, not the equation.

OpenStudy (anonymous):

lol never mind then im lost thanks tho ;)

OpenStudy (anonymous):

slope intercept form is y = 3x-4

OpenStudy (katrinakaif):

Haha, Well then, I aint leaving till you get this down. You found your slope which is 3. We need to now find the slope-intercept form and to do so, we use the formula: y-y1 = m (x-x1) Lets first substitute "m" or slope to be 3 y-y1 = 3 (x-x1) Now lets pick any one of the two pairs of given points: (-1,-7) OR (1,-1) I pick the second one, (1,-1) which we substitute in as (x1,y1) Now plug it into the equation we had earlier: y-y1 = 3 (x-x1) y - (-1) = 3 (x - 1) y + 1 = 3(x-1) Distribute the 3 to (x-1) y + 1 = 3x - 3 Bring the one to the other side by substracting it: y = 3x - 3 -1 y = 3x -4 Now THAT is your slope intercept form.

OpenStudy (anonymous):

yeah thanks and the standard form is y = 3x-4 -3x + y = -4 3x - y = 4 :)

OpenStudy (katrinakaif):

Lol Yep :] *Claps*

OpenStudy (anonymous):

thanks for your help precciate it